- 183: Sonics, Site stn6, Wed 27-Oct-1999 13:59:30 CDT, Sonic problem at Station3
The UW sonic anemometer v and w were bad for station 3.
At ~1200 all the connectors were checked and the transducers cleaned and
jiggled slightly.
No change was observed.
- 159: Sonics, Site stn6, Fri 22-Oct-1999 11:47:23 CDT, Replace ATI At Station 6
Noted yesterday the V component of the ATI at station 6 had a positive offset compared to rest of the sonics. Presious status has noted peculiarities in the response of this sonic. Replaced with the sensor brought out by TOny and CAthy on the 20th. Old sensor # is ATI 980302 (ATI#2). New Sensor ATI 98303.
- 142: Kryptons, Site stn6, Tue 19-Oct-1999 19:11:56 CDT, Remounted KH2O sensor on Station 6
Replaced the KH2O sensor on Station 6. This had been mounted at 20m
on the 55m tower for several days. We lowered the mast around 1835 CDT
and raised it again around 1900 CDT, Oct 19.
This means that the h2o.k.20m on the tower no longer exist!
However, it is still in the config files for ops3.
- 129: Sonics, Site stn6, Sun 17-Oct-1999 19:28:35 CDT, Sonic put back up at pam 6
Tom installed the sonic back on station 6. The mast was back in its
upright position at 17:15 CDT, Oct 17.
- 99: Kryptons, Site stn6, Thu 14-Oct-1999 18:07:39 CDT, Moved KH2O from Station 6 to 20m on 60m tower
Removed the KH2O from Station 6 at 15:30 local time, Oct 14, and
installed at 20m because we moved the Station 6 sonic to 20m when
the latter died this moring. Calibrated data not available until
data system was reconfigured at 17:30. It was assumed from logbook
entry #19 that this was S/N 1393. From the calibration sheet in
the case, we used Kw = -0.145, Intercept = 4388, Path = 1.29.